after scoring a strike, your bowling ball ( M = 7.2 kg, R = 11.0 cm) rolls, without slipping, back toward the ball rack, To reach the rack, the ball rolls up a ramp that rises through a vertical distance of 0.55 m. if the ball approaches the ramp with a speed of 2.800 m/s, what will be the speed of the ball, when it reaches the top of the ramp?
Assume that the ball is a solid sphere
Answer #1

solidsphere = 1 - 4 MR2 UE of rolling solid ophere = 2 mv 2 + 1 I wa NawR] = { mu2+ (fur?)(*) -7 ادم Vi = 2.8ms. Th=0.55m or